# Bioverse: GMT and ELT Direct Imaging and High-Resolution Spectroscopy Assessment – Surveying Exo-Earth O2 and Testing the Habitable Zone Oxygen Hypothesis

### v1

@@ Abstract
Biosignature detection in the atmospheres of Earth-like exoplanets is one of the most significant and ambitious goals for astronomy, astrobiology, and humanity. Molecular oxygen is among the strongest indicators of life on Earth, but it will be extremely difficult to detect via transmission spectroscopy. We used the Bioverse statistical framework to assess the ability to probe Earth-like O2 levels on hypothetical nearby habitable zone exoplanets (EECs) using direct imaging and high-resolution spectroscopy on the Giant Magellan Telescope (GMT) and the Extremely Large Telescope (ELT). We found that O2 could be probed on up to $\sim$5 and $\sim$15 EECs orbiting bright M dwarfs within 20 pc in a 10-year survey on the GMT and ELT, respectively. Earth-like O2 levels could be probed on four known super-Earth candidates, including Proxima Centauri b, within about one week on the ELT and a few months on the GMT. We also assessed the ability of the ELT to test the habitable zone oxygen hypothesis – that habitable zone Earth-sized planets are more likely to have O2 – within a 10-year survey using Bioverse. Testing this hypothesis requires either $\sim$1/2 of the EECs to have O2 or $\sim$1/3 if $\eta_{\oplus}$ is large. A northern hemisphere large-aperture telescope, such as the Thirty Meter Telescope (TMT), would expand the target star pool by about 25%, reduce the time to probe biosignatures on individual targets, and provide an additional independent check on potential biosignature detections.
@@ 1.1 Previous Direct Imaging Studies
Ground-based reflected-light spectroscopy of biosignatures necessitates the use of high resolution spectrographs to mitigate some of the effects of line blending with Earth’s atmosphere. Sparks & Ford (2002) proposed combining a coronographic imager to a high spectral resolution integral-field spectrograph for detection and characterization of exoplanets. As Snellen et al. (2015) showed, the achievable contrast of a combined high-contrast imaging and high-resolution spectrograph system is the product of the individual system achievable contrasts. For example, if each individual system can reach a contrast of $\sim$10-5, the combined system could reach contrasts of $\sim$10-10. Snellen et al. (2015) investigated using an optical high-resolution spectrograph (R=100,000) in combination with an extreme adaptive optics (AO) system on the ELT to observe a hypothetical 1.5 $R_{\oplus}$ planet in the habitable zone of Proxima Centauri in reflected light (prior to the discovery of Proxima Centauri b by Anglada-Escudé et al., 2016). They concluded that such a planet would be detectable in broadband (0.6–0.9 $\mu$m) reflected light in 10 hours of integration time at a S/N of $\sim$10 after cross-correlating the hypothetical observed spectrum with a model template spectrum, but they did not explore specific biosignature detection.
Since its discovery, most of the focus on high-contrast imaging and high-resolution reflected-light spectroscopy has been on Proxima Centauri b. This is likely driven by the planet’s location within the habitable zone, and a measured minimum mass ($M\sin i$) between $1.0~{}M_{\oplus}$ (Damasso et al., 2020) and $1.27~{}M_{\oplus}$ (Anglada-Escudé et al., 2016). Bixel & Apai (2017) placed probabilistic mass and radius constraints on Proxima Centauri b at $M=1.63^{+1.66}_{-0.72}~{}M_{\oplus}$ and $R=1.07^{+0.38}_{-0.31}~{}R_{\oplus}$, which is consistent with a rocky composition, but there is still a $\sim$10chance the planet has a significant amount of ice or a volatile envelope. We do not yet have enough information to truly constrain the nature of Proxima Centauri b, so assuming it is Earth-like or an Earth-analog is an optimistic assumption.
@@
Wang et al. (2017) simulated observations of Proxima Centauri b (assuming it is an Earth-like planet) and a hypothetical Earth-like habitable zone planet orbiting an M dwarf at 5 pc using a 30-meter class telescope at near-infrared bands. They used near-infrared (1–2.5 $\mu$m) R=1,000–100,000 Earth-like exoplanet spectra, considering atmospheric chemistry and using a radiative transfer model from Hu et al. (2012a, b, 2013); Hu & Seager (2014). For a 100-hour integration time, their simulations yielded S/N$>$10 for H2O, O2, CO2, and CH4, typically when R$>$10,000 and star-light suppression was $>10^{-7}$. These simulations simplistically addressed line blending, assuming the radial velocity of the exoplanet shifts its atmospheric signature by tens of km s-1. An updated version of this analysis was provided by Zhang et al. (2024), where they simulated direct imaging of 10 nearby rocky planets at R=1000 with HARMONI and METIS on the ELT in search of biosignatures. They indicated that CO2, CH4, and H2O should be detectable on GJ 887 b and Proxima Centauri b at a S/N$>$5 with integration times on the order of $<$100 hours.
@@ 1.2 Bioverse
Bioverse was developed by Bixel & Apai (2021) as a modular framework to generate and classify exoplanets based on exoplanet demographics, simulate exoplanet surveys for upcoming telescopes and missions, and compute the diagnostic power of future surveys in testing population-level hypotheses. Bioverse folds in state-of-the-art exoplanet occurrence rate calculations, incorporates a complete nearby star catalog out to $\sim$100 pc, and accounts for realistic ground-based observing constraints, allowing exploration of the full potential of a broad statistical survey of nearby exoplanets.
@@
After introducing the framework in Bixel & Apai (2021), Hardegree-Ullman et al. (2023) added a stellar catalog based on Gaia Data Release 3, the capability to determine transit observability for ground-based large-aperture telescopes, and simulations for oxygen detection via transmission spectroscopy. Schlecker et al. (2024) introduced selectable mass-radius relations, a mission simulator for ESA’s PLATO mission (Rauer et al., 2016), and a runaway greenhouse climate model. Here, we built upon these previous studies to assess the direct imaging and high resolution spectroscopy potential for reflected light studies with the GMT and ELT.
@@ 2.1 Planet Generation and Observational Constraints
In order to provide a realistic survey simulation, we used Bioverse to generate a set of exoplanets orbiting bright stars (Gaia red-band $G_{RP}\leq 10$) out to 20 pc from the Sun based on planet occurrence rates from Bergsten et al. (2022), which was first introduced to Bioverse in Schlecker et al. (2024). The underlying parametric model in Bergsten et al. (2022) is derived from planet radius, orbital period, and stellar host mass distributions in Kepler’s FGK and early M stars sample, considering planets with radii up to $3.5~{}R_{\oplus}$ and orbiting out to 100 days. The original SAG 13-based (Kopparapu et al., 2018) planet occurrence rate module introduced by Bixel & Apai (2021) was derived from a compilation of Kepler planet occurrence rates prior to 2017, and did not include a stellar mass dependence. Both of the planet occurrence rate modules are model-dependent for planets orbiting beyond 100 days.
@@
$\mathrm{RV}_{\mathrm{orb}}=\sqrt{GM_{\star}\left(\frac{2}{r}-\frac{1}{a}\right)}$ (2)
where $G$ is the gravitational constant, $M_{\star}$ is the mass of the host star, $r$ is the distance of the exoplanet to the host star at the time of observation, and $a$ is the semi-major axis of the exoplanet. Observer barycentric velocity at the time of observation was computed using the baryCorr function from the PyAstronomy package (Czesla et al., 2019). For an $R=100,000$ spectrograph observing the O2 A-band, severe line blending occurs when $|\Delta\mathrm{RV}|<13$ km s-1 and $30\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<55\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023).
@@
$\frac{F_{p}(\lambda,\alpha)}{F_{\star}(\lambda)}=A_{g}(\lambda)g(\alpha)\left(\frac{R_{p}}{a}\right)^{2},$ (3)
@@
where the illumination phase angle $\alpha=\cos^{-1}(-\sin i\cos\phi)$, $i$ is the exoplanet inclination angle, and $\phi$ is the orbital phase. Geometric albedo, inclination angle, and orbital phase were randomly generated from uniform distributions for each simulated planet within Bioverse.
@@ 2.2 Instrument Models
In order to provide a realistic simulation of extreme-AO performance, we need models for planned instruments, which currently include GMagAO-X (Males et al., 2022) for the GMT and the Planetary Camera and Spectrograph (PCS, Kasper et al., 2021) for the ELT. However, full end-to-end simulations of these high-order AO systems are very expensive in terms of computational time and complexity. Therefore, we used a semi-analytical approach to the performance simulation (Jolissaint, 2010). The semi-analytical approach assumes that all temporal evolution of the turbulence is driven by Taylor’s frozen flow. The frozen flow approximates the short time evolution by shifting the spatial structure of the turbulence with single wind velocity. The time evolution of any parameter can then be replaced by a shifted version, $\varphi(\vec{r},t)=\varphi(\vec{r}-\vec{v}t)$. This time evolution underpins the semi-analytical model and allows us to express an adaptive optics system as a series of spatial filters that act on the turbulence Power Spectral Density (PSD). The AO system is controlled using integral control, $\mathrm{DM}_{i+1}=\mathrm{DM}_{i}-g\varepsilon_{i}$. Here, $\varepsilon_{i}$ is the wavefront error at time step $i$ and $\mathrm{DM}_{i}$ is the DM shape at time step $i$. The feedback gain $g$ is a function of spatial frequency and was optimized following Gendron & Léna (1994); Jolissaint (2010).
@@
We simulated two different representative AO system optimizations. In the first, the AO system was allowed to run as fast as possible, which is 3 kHz for both GMagAO-X and PCS. The gain for each spatial frequency was optimized at this AO loop speed. In the second approach, we optimized the Strehl by changing both the AO loop speed and modal gains. These two approaches were used because optimizing for just Strehl does not necessarily lead to the best contrast. The contrast at small angular separations is driven by the servo lag error. This error term is a larger contributor on GMT/ELT-sized instruments than on current instruments (Males & Guyon, 2018). Therefore, it is possible to increase the contrast at smaller angular separations by running the AO system faster at the cost of lower Strehl. This leads to a better final signal-to-noise. The two different approaches are called ‘Max Speed’ and ‘Strehl Optimization’ in this work. A more thorough assessment of which instrument and observing modes were selected for scientific analysis in Section 3 is given in Appendix A.
@@
The AO performance depends strongly on the assumed atmospheric conditions. For our simulations, we followed Prieto et al. (2010) and Thomas-Osip et al. (2010) for the GMT. The first quartile conditions are $r_{0}=0.22$ m and a $C_{n}^{2}$ integrated wind velocity of 9.4 m s-1. The second quartile conditions are $r_{0}=0.17$ m and a mean wind speed of 18.4 m s-1. For the ELT, we followed Sarazin et al. (2013): the first and second quartile conditions are $r_{0}=0.24$ m with a velocity of $10$ m s-1 and $r_{0}=0.15$ m with a velocity of 11 m s-1, respectively.
@@
The design of GMagAO-X is currently further along than that of the other extreme-AO systems (Males et al., 2022; Haffert et al., 2022). Therefore, we used the current design parameters for GMagAO-X as reference. Specifically, we assumed 182 actuators across the pupil and a pyramid wavefront sensor. The design of PCS is currently still in its conceptual phase and no hard decisions have been made (Kasper et al., 2021). Therefore, we considered two different DM sizes for PCS: 200 actuators across the pupil based on the proposed density for EPICS (Kasper et al., 2010), and 128 actuators across the pupil based on a more recent ELT extreme-AO design (Madec et al., 2022).
@@
The output of the semi-analytical model is a post-AO PSD. The PSD still has to be propagated through an actual coronagraphic system to measure the amount of residual scattered star light. The expected exposure times of our simulations are on the order of hours. This is much longer than the typical speckle lifetime of 1 to 10 ms (Males et al., 2021). We can treat the different spatial frequencies as incoherent because of the many orders of magnitude differences in characteristic timescales. With that in mind, the propagation through a coronagraph can be viewed as an incoherent sum of plane waves (Sauvage et al., 2010; Herscovici-Schiller et al., 2017). Finally, the effects of quasi-static speckles are ignored in the simulations for two reasons. First, for high-resolution spectroscopy the speckle lifetime does not matter, only the amount of residual starlight matters. This is not true for other post-processing techniques such as Angular Differential Imaging, which is extremely sensitive to the speckle lifetime. Second, controlling and removing quasi-static speckles below the residual atmospheric speckles is now possible (Potier et al., 2022; Haffert et al., 2023; Galicher et al., 2024). Therefore, no quasi-static speckles were added to our simulations, under the assumption that the efficiency of quasi-static speckle removal process will continue to improve over the next few years.

### v2

@@ Abstract
Biosignature detection in the atmospheres of Earth-like exoplanets is one of the most significant and ambitious goals for astronomy, astrobiology, and humanity. Molecular oxygen is among the strongest indicators of life on Earth, but it will be extremely difficult to detect via transmission spectroscopy. We used the Bioverse statistical framework to assess the ability to probe Earth-like O2 levels on hypothetical nearby habitable zone exo-Earth candidates (EECs) using direct imaging and high-resolution spectroscopy on the Giant Magellan Telescope (GMT) and the Extremely Large Telescope (ELT). Assuming continued improvement in instruments and data processing, our analysis highlights the best-case scenarios. Earth-like O2 levels could be probed on up to $\sim$7 and $\sim$19 EECs orbiting bright M dwarfs within 20 pc in a hypothetical 10-year survey on the GMT and ELT, respectively. Four known super-Earth candidates, including Proxima Centauri b, could be probed for O2 within about one week of observations on the ELT and a few months on the GMT. We also assessed the ability of the ELT to test the habitable zone oxygen hypothesis – that habitable zone Earth-sized planets are more likely to have O2 – within a 10-year survey using Bioverse. Testing this hypothesis requires either $\sim$1/2 of the EECs to have O2 or $\sim$1/3 if $\eta_{\oplus}$ is large. A northern hemisphere large-aperture telescope, such as the Thirty Meter Telescope (TMT), would expand the target star pool by about 25%, reduce the time to probe biosignatures on individual targets, and provide an additional independent check on potential biosignature detections.
@@ 1.1 Previous Direct Imaging Studies
Ground-based spectroscopy of biosignatures necessitates the use of high resolution spectrographs ($R\gtrsim 50,000$) to allow separation of Earth’s telluric lines from those of the exoplanet atmosphere, assuming sufficient line Doppler shifts to minimize line blending (see, e.g., Rodler & López-Morales, 2014; López-Morales et al., 2019, for examples of different resolutions and their effect on line blending). Sparks & Ford (2002) proposed combining a coronographic imager to a high spectral resolution integral-field spectrograph for detection and characterization of exoplanets. As Snellen et al. (2015) showed, the achievable contrast of a combined high-contrast imaging and high-resolution spectrograph system is the product of the individual system achievable contrasts. For example, if each individual system can reach a contrast of $\sim$10-5, the combined system could reach contrasts of $\sim$10-10 (see additional discussion in Section 2.2).
@@
Snellen et al. (2015) investigated using an optical high-resolution spectrograph (R=100,000) in combination with an extreme adaptive optics (AO) system on the ELT to observe a hypothetical 1.5 $R_{\oplus}$ planet in the habitable zone of Proxima Centauri in reflected light (prior to the discovery of Proxima Centauri b by Anglada-Escudé et al., 2016). They concluded that such a planet would be detectable in broadband (0.6–0.9 $\mu$m) reflected light in 10 hours of integration time at a S/N of $\sim$10 after cross-correlating the hypothetical observed spectrum with a model template spectrum, but they did not explore specific biosignature detection.
Since its discovery, most of the focus on high-contrast imaging and high-resolution reflected-light spectroscopy has been on Proxima Centauri b. This is likely driven by the planet’s location within the habitable zone, and a measured minimum mass ($M\sin i$) likely between $1.0~{}M_{\oplus}$ (Damasso et al., 2020) and $1.27~{}M_{\oplus}$ (Anglada-Escudé et al., 2016). Bixel & Apai (2017) placed probabilistic mass and radius constraints on Proxima Centauri b at $M=1.63^{+1.66}_{-0.72}~{}M_{\oplus}$ and $R=1.07^{+0.38}_{-0.31}~{}R_{\oplus}$, which is consistent with a rocky composition, but there is still a $\sim$10chance the planet has a significant amount of ice or a volatile envelope. We do not yet have enough information to truly constrain the nature of Proxima Centauri b, so assuming it is Earth-like or an Earth-analog is an optimistic assumption.
@@
Wang et al. (2017) simulated observations in reflected light with low to high-resolution spectra of Proxima Centauri b (assuming it is an Earth-like planet) and a hypothetical Earth-like habitable zone planet orbiting an M dwarf at 5 pc using a 30-meter class telescope at near-infrared bands. They used near-infrared (1–2.5 $\mu$m) R=1,000–100,000 Earth-like exoplanet spectra, considering atmospheric chemistry and using a radiative transfer model from Hu et al. (2012a, b, 2013); Hu & Seager (2014). For a 100-hour integration time, their simulations yielded S/N$>$10 for H2O, O2, CO2, and CH4, typically when R$>$10,000 and star-light suppression was $>10^{-7}$. These simulations simplistically addressed line blending, assuming the radial velocity of the exoplanet shifts its atmospheric signature by tens of km s-1. An updated version of this analysis was provided by Zhang et al. (2024), where they simulated direct imaging of 10 nearby rocky planets at R=1000 with HARMONI-like and METIS-like instruments on the ELT in search of biosignatures. They indicated that CO2, CH4, and H2O should be detectable on GJ 887 b and Proxima Centauri b at a S/N$>$5 with integration times on the order of $<$100 hours.
Hawker & Parry (2019) focused on the detectability of Earth-like levels of O2 on Proxima Centauri b with high-resolution spectroscopy and high contrast imaging on the ELT using HIRES and HARMONI-like instruments. They concluded a S/N=3–5 detection could be obtained in 30–70 hours of integration time. More recently, Vaughan et al. (2024) simulated observations of Proxima Centauri b using HARMONI on the ELT at its highest resolving power of R=17,385. Their simulations indicated an atmosphere on Proxima Centauri b (assuming it is Earth-like) could be characterized at S/N=5 in 20–30 hours of integration time. These observations would be particularly sensitive to CH4, but CO2 would be harder to detect.
@@ 1.2 Bioverse
Bioverse was developed by Bixel & Apai (2021) as a modular framework to generate and classify exoplanets based on exoplanet demographics, simulate exoplanet surveys for upcoming telescopes and missions, and compute the diagnostic power of future surveys in testing population-level hypotheses. Bioverse folds in state-of-the-art exoplanet occurrence rate calculations, incorporates a complete nearby star catalog out to $\sim$100 pc (Gaia red-band magnitude range $19.90.5<G_{RP}<$ 19.9), and accounts for realistic ground-based observing constraints, allowing exploration of the full potential of a broad statistical survey of nearby exoplanets.
@@
After introducing the framework in Bixel & Apai (2021), Hardegree-Ullman et al. (2023) added a stellar catalog based on Gaia Data Release 3, the capability to determine transit observability for ground-based large-aperture telescopes, and simulations for oxygen detection via transmission spectroscopy. Schlecker et al. (2024) introduced selectable mass-radius relations, a mission simulator for ESA’s PLATO mission (Rauer et al., 2016), and a runaway greenhouse climate model. Here, we built upon these previous studies to assess the direct imaging and high resolution spectroscopy potential for reflected light studies with the GMT and ELT, specifically focusing on the ability to probe Earth-like O2 levels. Leveraging all modules of Bioverse, we seek to test the “habitable zone oxygen hypothesis,” i.e., the hypothesis that Earth-sized planets within a star’s habitable zone are more likely to have Earth-like O2 levels. This will inform us if a hypothetical 10-year survey of EECs with the ELT will be sufficient to test this hypothesis.
@@ 2.1 Planet Generation and Observational Constraints
In order to provide a realistic survey simulation, we used Bioverse to generate a set of exoplanets orbiting bright stars (Gaia red-band $G_{RP}\leq 10$) out to 20 pc from the Sun based on planet occurrence rates from Bergsten et al. (2022), which was first introduced to Bioverse in Schlecker et al. (2024). The underlying parametric model in Bergsten et al. (2022) is derived from planet radius, orbital period, and stellar host mass distributions in Kepler’s FGK and early M stars sample, considering planets with radii up to $3.5~{}R_{\oplus}$ and orbiting out to 100 days. The original SAG 13-based (Kopparapu et al., 2018) planet occurrence rate module introduced by Bixel & Apai (2021) was derived from a compilation of Kepler planet occurrence rates prior to 2017, and did not include a stellar mass dependence. Both of the planet occurrence rate modules are model-dependent for planets orbiting beyond 100 days. This does not affect our study much because an overwhelming majority of targets in our simulated surveys are M dwarfs (see Figure 5) with habitable zones closer than 100 days. We note that the SAG 13 occurrence rates are more optimistic than the Bergsten et al. (2022) values by a factor of $\sim$2, but we opt to use the Bergsten et al. (2022) values in this study because it is not only more recent, but also from a uniform study rather than a compilation.
@@
$\mathrm{RV}_{\mathrm{orb}}=\sqrt{GM_{\star}\left(\frac{2}{r}-\frac{1}{a}\right)},$ (2)
where $G$ is the gravitational constant, $M_{\star}$ is the mass of the host star, $r$ is the distance of the exoplanet to the host star at the time of observation (computed in Bioverse using Newton’s method to numerically solve Kepler’s equation), and $a$ is the semi-major axis of the exoplanet. Observer barycentric velocity at the time of observation was computed using the baryCorr function from the PyAstronomy package (Czesla et al., 2019). For an $R=100,000$ spectrograph observing the O2 A-band, severe line blending occurs when $|\Delta\mathrm{RV}|<13$ km s-1 and $30\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<55\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023).
@@
$C=\frac{F_{p}(\lambda,\alpha)}{F_{\star}(\lambda)}=A_{g}(\lambda)g(\alpha)\left(\frac{R_{p}}{a}\right)^{2},$ (3)
@@
[…]
@@ 2.2 Instrument Models
In order to provide a realistic simulation of extreme-AO performance, we need models for planned instruments, which currently include GMagAO-X (Males et al., 2022) for the GMT and the Planetary Camera and Spectrograph (PCS, Kasper et al., 2021) for the ELT. However, full end-to-end simulations of these high-order AO systems are very expensive in terms of computational time and complexity. Therefore, we used a semi-analytical approach to the performance simulation (Jolissaint, 2010). The semi-analytical approach assumes that all temporal evolution of the turbulence is driven by Taylor’s frozen flow. The frozen flow approximates the short time evolution by shifting the spatial structure of the turbulence with single-wind velocity. The time evolution of any parameter, $f$ for example, can then be replaced by a shifted version, $f(\vec{r},t)=f(\vec{r}-\vec{v}t)$. This time evolution underpins the semi-analytical model and allows us to express an adaptive optics system as a series of spatial filters that act on the turbulence Power Spectral Density (PSD). The spatial frequency approach has been verified extensively in comparison with full end-to-end models (Males et al., 2021). The AO system is controlled using integral control, $\mathrm{DM}_{i+1}=\mathrm{DM}_{i}-g\varepsilon_{i}$. Here, $\varepsilon_{i}$ is the wavefront error at time step $i$ and $\mathrm{DM}_{i}$ is the DM shape at time step $i$. The feedback gain $g$ is a function of spatial frequency and was optimized following Gendron & Léna (1994); Jolissaint (2010).
@@
We simulated two different representative optimizations of the AO system. In the first, the AO system was allowed to run as fast as possible, which is 3 kHz for both GMagAO-X and PCS. The gain for each spatial frequency was optimized at this AO loop speed. In the second approach, we optimized the Strehl by changing both the AO loop speed and modal gains. These two approaches were used because optimizing for only Strehl does not necessarily lead to the best contrast. The contrast at small angular separations is driven by the servo lag error. This error term is a larger contributor on GMT/ELT-sized instruments than on current instruments (Males & Guyon, 2018). Therefore, it is possible to increase the contrast at smaller angular separations by running the AO system faster at the cost of lower Strehl. This leads to a better final signal-to-noise ratio. The two different approaches are called ‘Max Speed’ and ‘Strehl Optimization’ in this work. A more thorough assessment of which instrument and observing modes were selected for scientific analysis in Section 3 is given in Appendix A.
@@
AO performance strongly depends on the atmospheric conditions assumed. For our simulations, we followed Prieto et al. (2010) and Thomas-Osip et al. (2010) for the GMT (see Table 1). The first quartile conditions have a Fried parameter of $r_{0}=0.22$ m and a $C_{n}^{2}$ integrated wind velocity of 9.4 m s-1. Here $C_{n}^{2}$ is the refractive index structure parameter. The second quartile conditions are $r_{0}=0.17$ m and a mean wind speed of 18.4 m s-1. For the ELT, we followed Sarazin et al. (2013): the first and second quartile conditions are $r_{0}=0.24$ m with a velocity of $10$ m s-1 and $r_{0}=0.15$ m with a velocity of 11 m s-1, respectively.
@@
The design of GMagAO-X is currently further along than that of the other extreme-AO systems (Males et al., 2024; Haffert et al., 2024). Therefore, we used the current design parameters for GMagAO-X as reference. Specifically, we assumed 188 actuators across the pupil and a pyramid wavefront sensor. The design of PCS is currently still in its conceptual phase and no hard decisions have been made (Kasper et al., 2021). Therefore, we considered two different DM sizes for PCS: 200 actuators across the pupil based on the proposed density for EPICS (Kasper et al., 2010), and 128 actuators across the pupil based on a more recent ELT extreme-AO design (Madec et al., 2022).
@@
The output of the semi-analytical model is a post-AO PSD. The PSD still has to be propagated through an actual coronagraphic system to measure the amount of residual scattered star light. The expected exposure times of our simulations are on the order of hours. This is much longer than the typical speckle lifetime of 1 to 10 ms (Males et al., 2021). We can treat the different spatial frequencies as incoherent because of the many orders of magnitude differences in characteristic timescales. With that in mind, the propagation through a coronagraph can be viewed as an incoherent sum of plane waves (Sauvage et al., 2010; Herscovici-Schiller et al., 2017). The theoretical perfect coronagraph removes only the piston mode from the incoming wavefront (Guyon et al., 2006; Deshler et al., 2024). However, such coronagraphs are extremely sensitive to residual jitter and partially-resolved stars (Guyon et al., 2006). Therefore, we used a 4th-order perfect coronagraph that also removes the tip/tilt modes from the wavefronts beside the piston mode. This increases the inner-working angle of the coronagraph and it increases the system robustness against stellar angular diameter and residual tip/tilt jitter(Guyon et al., 2006; Belikov et al., 2021). Given the size of the GMT/ELT, it will not be realistic to work with a 2nd-order coronagraph that only removes pistons. Finally, the effects of time varying quasi-static speckles are ignored in the simulations for two reasons. For high-resolution spectroscopy, the only thing that matters is the amount of starlight that is left in the final spectra. The cross-correlation function is differentiating between stellar residuals and planet light through spectral information only and the associated noise is photon noise of the time-integrated spectrum intensity (Snellen et al., 2015). The speckle lifetime affects the variance of this intensity but not the mean (shorter lifetime $\rightarrow$ smaller variance/mean). We assume that the quasi-static speckles can be controlled and removed to a level that is below the halo of residual atmospheric speckles using newly implemented focal plane wavefront control techniques (Potier et al., 2022; Haffert et al., 2023; Galicher et al., 2024) and that means the photon noise from the atmospheric speckles will dominate the noise budget. Therefore, the total power of the low-spatial frequency content of the quasi-static speckles was set to 1 nm and the high-spatial frequency content was set to 25 nm and added to our simulations. This amount is similar to what has been achieved with MagAO-X (Van Gorkom et al., 2021). We also assume that the efficiency of quasi-static speckle removal process will continue to improve over the next few years. An example output of the semi-analytical modeling is shown in Figure 1.
Figure 1: Output of the semi-analytical AO simulations for ELT/PCS with the 128 actuators model. The left figure shows the simulated PSF and the right figure shows the coronagraphic stellar image. There are two visible circular regions. The largest corresponds to the control radius of the AO system and the smaller one corresponds to the control radius of the non-common path aberrations. The contrast within the inner radius is dominated by the wind-driven halo. Axes are shown in both mas and $\lambda/D$, assuming $\lambda=765$ nm and $D=39.3$ m.
Figure 2: Coronagraphic encircled energy throughput ($1.8\lambda/D$ diameter) as a function of angular separation for an 8th magnitude star. A 4th-order perfect coronagraph with the ELT aperture was used in these simulations. The top row shows the PSFs corresponding to the four orange points on the throughput curve from 0 mas (left) to 20 mas (right) separation. The PSF is deformed close to the inner-working angle of the coronagraph.
ELT Parameters
Telescope area 978 m2
Number of actuators 128 / 200
Number of wfs pixels 125,000
GMT Parameters
Telescope area 358 m2
Number of actuators 188
Number of wfs pixels 111,000
Shared AO System Parameters
Zeropoint flux in I band 4520 photons s-1 cm-2 nm-1
Bandwidth I band 150 nm
WFS wavelength 765 nm
Loop speed 250 - 4000 Hz
Intrinsic delay 1.5 frames
Computational delay 250 $\mu$s
Throughput to WFS camera 10 %
WFS detector variance 2.0 $e^{-}$
Photon noise sensitivity $1/\sqrt{2}$
Read noise sensitivity $1/\sqrt{2}$
Common path throughput to spectrograph 25 %
Coronagraph 4th order perfect coronagraph (Guyon et al., 2006)
Site Parameters
$r_{0}$ 10 - 20 cm
$L_{0}$ 25 m
$v$ 15 - 40 m s-1
Spectrograph Parameters
read noise $1e^{-}$ pixel-1
dark current $10^{-4}$ $e^{-}$ s-1 pixel-1
Spectrograph throughput 35 %
Detector quantum efficiency 86 %
Science wavelength 765 nm
Spectral bandwidth 10 nm
Resolving power 100,000 – 500,000
Spectral resolution 3 channels per resolving element
Pixels per channel 2 pixels
Table 1: Parameters used to calculate S/N (Equation 5).
In order to calculate the signal-to-noise ratio of the high-contrast, high resolution spectroscopy observations of the exoplanet atmosphere O2 A-band, we follow the basic principle of Snellen et al. (2015) and Lovis et al. (2017):
{widetext}
$S/N=\frac{\eta\Delta tT_{p}T_{I}T_{AO}T_{QE}CF_{\star}}{\sqrt{\Delta tT_{\star}T_{I}T_{AO}T_{QE}F_{\star}+\Delta t\sigma^{2}_{\mathrm{sky}}+N_{\mathrm{exp}}N_{\mathrm{pixels}}\sigma^{2}_{\mathrm{read}}+N_{\mathrm{pixels}}\Delta t\sigma^{2}_{\mathrm{dark}}}},$ (5)
[…]
At high resolution ($R\geq 100,000$), we can mitigate Earth’s telluric signal by avoiding observing the exoplanet at relative radial velocities at which line blending is severe (Section 2.1). Disentangling the star and exoplanet signal happens in post-processing by cross-correlating the reflected-light spectrum with the stellar spectrum. Cross-correlation post-processing efficiency ($\eta$) for the O2 A-band converges to 95% for $R\geq 100,000$ (Fowler et al., 2023) if the planet is sufficiently Doppler-shifted (Hardegree-Ullman et al., 2023). Rearranging Equation 5, we computed the achievable contrast $C$ at S/N=5 for 1-hour total integration times (assuming four 15-minute exposures) for different host star apparent magnitudes and angular separations for the GMT (Table 3) and ELT (Table 4). The brightest stars might require shorter exposures, but the dominant detector noise will be dark current, which scales with exposure time, rather than read noise, so the effect is insignificant.
Figure 3: The S/N after a 1 hour integration for different planet contrasts and stellar magnitudes. The S/N=5 detection limit is shown by the dashed grey line.
[…]